OpenStudy (anonymous):

46 is what percent of 107

OpenStudy (ja1):

\[Percent = \frac{ 46 }{ 107 }\]

OpenStudy (anonymous):

Can you walk me through it please

OpenStudy (ja1):

Of course! Basically 46/107 means we need to divide 46 by 107, so we can think of it like this: \[46 \div 107\]

OpenStudy (ja1):

So for this we can use a calculator to divide that, do you have one handy?

OpenStudy (anonymous):

0.4299065420560748

OpenStudy (anonymous):

Yes I do

OpenStudy (anonymous):

Thats Whatcha Get Out Of The Calculator

OpenStudy (ja1):

Well oliver already did the calculation for us :) So now we need to take that number and round it. (If that is what your instructor wants.)

OpenStudy (anonymous):

No Problem You Two :)

OpenStudy (anonymous):

No he doesn't want us rounding

OpenStudy (ja1):

Ok, then you basically already have the answer there :)

OpenStudy (anonymous):

0.42

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Thank @JA1

OpenStudy (mathstudent55):

English: 46 is what percent of 107 Math: 46 = x % * 107 46 = (x/100) * 107 46 = 107x/100 46*100 = 107x 46*100/107 = x